Which of the following odd numbers can you write as the sum of two prime numbers? Try to write them as sum of three odd numbers.
1. 35
2. 31
3. 65

Answer by ikleyn(53763)

.

First question


    31 = 2 + 29.   This and only this number.


    HINT.  One of the two addends must be 2, the only even prime number.


Second question

    35 = 3 + 13 + 19  (the sum of three odd prime numbers)


    31 = 3 + 11 + 17  (the sum of three odd prime numbers)


    65 = 3 + 19 + 43  (the sum of three odd prime numbers)

Solved.